| #1423 |
02 Apr 08:25 Signal Chain stop
|
new
|
J. Ortiz
|
Process/Analysis Program
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1424 |
02 Apr 10:01 Signal Chain restart
|
new
|
J. Ortiz
|
Process/Analysis Program
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1425 |
02 Apr 12:01 Drop in Tx line
|
new
|
J. Ortiz
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1426 |
02 Apr 12:22 Calibration of ATRs and TRs
|
new
|
J. Ortiz
|
Antennas
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1427 |
02 Apr 13:34 Posible Solar Flare Interference
|
new
|
J. Ortiz
|
Physical Phenomena
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1428 |
02 Apr 14:40 Power adjustment in Tx1 due to AC Line variation
|
new
|
J. Ortiz
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1429 |
02 Apr 15:20 Signal Chain Restart
|
new
|
J. Ortiz
|
Process/Analysis Program
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1430 |
02 Apr 15:22 Calibration of ATRs and TRs
|
new
|
J. Ortiz
|
Antennas
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1431 |
02 Apr 18:35 Calibration of ATRs and TRs
|
new
|
J. Ortiz
|
Antennas
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1432 |
02 Apr 19:50 Transmitter Power measurements
|
new
|
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1433 |
03 Apr 00:12 Transmitter Power measurements
|
new
|
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1434 |
02 Apr 21:55 Calibration of ATRs and TRs
|
new
|
jverastegui
|
Antennas
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1435 |
03 Apr 00:00 Figures of RTI, noise, spectras, etc
|
new
|
jverastegui
|
Process/Analysis Program
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1436 |
03 Apr 01:45 Calibration of ATRs and TRs
|
new
|
jverastegui
|
Antennas
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1437 |
03 Apr 07:10 Transmitter Power measurements
|
new
|
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1438 |
03 Apr 07:10 Transmitter Power measurements
|
new
|
|
Transmitters
|
normal
|
|
| #1439 |
03 Apr 07:05 Crowbar by voltage drop
|
new
|
jverastegui
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1440 |
02 Abr 07:50 Lose synchronization in the Faraday aquisition system
|
new
|
|
Acquisition system
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1441 |
03 Apr 08:41 Calibration of ATRs and TRs
|
new
|
dsuarez
|
General
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1442 |
03 Apr 11:53 Calibration of ATRs and TRs
|
new
|
dsuarez
|
General
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1443 |
03 Apr 15:09 Calibration of ATRs and TRs
|
new
|
dsuarez
|
General
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1444 |
03 Apr 17:03 IDL program is hanged
|
new
|
dsuarez
|
Process/Analysis Program
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1445 |
03 Apr 18:10 Calibration of ATRs and TRs
|
new
|
dsuarez
|
General
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1446 |
03 Apr 20:00 Transmitter Power mesearuments
|
new
|
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1447 |
03 Apr 22:20 spark gaps adjustment
|
new
|
I. Manay
|
Spark Gaps
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1448 |
04 Apr 23:50 Transmitter Power Measurements
|
new
|
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1449 |
03 Apr 23:30 Re-starting acquisition of the Faraday experiment
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1450 |
03 Apr 23:57 Re-starting acquisition of the Faraday experiment
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1451 |
03 Apr 23:59 plots of experiments
|
new
|
I. Manay
|
Plot
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1452 |
04 Apr 02:40 spark gaps adjustment
|
new
|
J. Lavado
|
Spark Gaps
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1453 |
04 Apr 06:55 spark gaps adjustment
|
new
|
J. Lavado
|
Spark Gaps
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1454 |
04 Apr 07:05 Transmitter Power Measurements
|
new
|
|
Transmitters
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1455 |
04 Apr 07:49 Re-starting acquisition of the Faraday experiment
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1456 |
04 Abr 12:30 Transmitter Power mesearuments
|
new
|
|
Transmitters
|
normal
|
2013_04_1 EW_Drift+Faraday+Imaging 29/04-03/05
|
| #1457 |
04 Abr 12:35 End of campaign
|
new
|
P. Condor
|
End experiment
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1458 |
09 Apr 13:30 Experiment data to PUMA
|
new
|
I. Manay
|
General
|
normal
|
2014_03 ISCOD HLStormtime
|
| #1459 |
21 Apr 14:19 Log created for 2014_04 ISR+Imaging
|
new
|
ryanque
|
Start experiment
|
normal
|
2014_04_ISR+Imaging
|
| #1460 |
21 Apr 02:22 Antenna VSWR
|
new
|
ryanque
|
Antennas
|
normal
|
2014_04_ISR+Imaging
|
| #1461 |
22 Apr 12:42 Starts campaign
|
new
|
ryanque
|
Start experiment
|
normal
|
2014_04_ISR+Imaging
|
| #1462 |
22 Apr 14:10 Drop at the channel A of Faraday experiment
|
new
|
imanay
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1463 |
22 Apr 15:00 spark gaps adjusment
|
new
|
imanay
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1464 |
22 Apr 15:55 Problems with Faraday acquisition system
|
new
|
imanay
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1465 |
23 Apr 08:35 Problems with Faraday acquisition system
|
new
|
J. Ortiz
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1466 |
23 Apr 08:51 Drop at the channel A of Faraday experiment
|
new
|
J. Ortiz
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1467 |
23 Apr 09:16 Drop in transmitter power source
|
new
|
J. Ortiz
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1468 |
23 Apr 09:22 Transmitter maintenance
|
new
|
J. Ortiz
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1469 |
23 Apr 11:48 Calibration of ATRs and TRs
|
new
|
|
General
|
normal
|
2014_04_ISR+Imaging
|
| #1470 |
23 Apr 11:50 Re-starting experiment
|
new
|
J. Ortiz
|
Start experiment
|
normal
|
2014_04_ISR+Imaging
|
| #1471 |
23 Apr 16:32 Experiment stop
|
new
|
J. Ortiz
|
End experiment
|
normal
|
2014_04_ISR+Imaging
|
| #1472 |
23 Apr 16:51 Imaging Experiment start
|
new
|
J. Ortiz
|
Start experiment
|
normal
|
2014_04_ISR+Imaging
|
| #1473 |
23 Apr 17:05 Calibration of ATRs and TRs
|
new
|
|
General
|
normal
|
2014_04_ISR+Imaging
|
| #1474 |
23 Apr 19:42 Unknown peak in data
|
new
|
J. Ortiz
|
Physical Phenomena
|
normal
|
2014_04_ISR+Imaging
|
| #1475 |
23 Apr 19:50 Transmitter Power Measurements
|
new
|
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1476 |
23 Abril 23:05 Transmitter Power Measurements
|
new
|
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1477 |
24 April 00:00 Driever 3 OFF
|
new
|
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1478 |
23 April 21:10 Calibration of ATRs and TRsF
|
new
|
AValdez
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1479 |
23 April 21:50 Problems in the monitoring program Imaging_plot_coherence.py
|
new
|
AValdez
|
Process/Analysis Program
|
normal
|
2014_04_ISR+Imaging
|
| #1480 |
23 April 23:55 SYNCHRONISM LOSS IN EW_DRIFT EXPERIMENT
|
new
|
AValdez
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1481 |
23 April 23:59 IDL program is hung EW_Drift
|
new
|
AValdez
|
Process/Analysis Program
|
normal
|
2014_04_ISR+Imaging
|
| #1482 |
24 April 00:20 Calibration of ATRs and TRs
|
new
|
AValdez
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1483 |
23 Apr 23:59 plots of experiments
|
new
|
AValdez
|
Plot
|
normal
|
2014_04_ISR+Imaging
|
| #1484 |
24 April 07:10 Transmitter Power Measurements
|
new
|
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1485 |
24 April 04:24 SYNCHRONISM LOSS IN EW_DRIFT EXPERIMENT
|
new
|
AValdez
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1486 |
24 April 04:24 SYNCHRONISM LOSS IN EW_DRIFT EXPERIMENT
|
new
|
AValdez
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1487 |
24 April 04:30 am Calibration of ATRs and TRs
|
new
|
AValdez
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1488 |
24 Apr 08:11 Calibration of ATR and TR
|
new
|
ryanque
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1489 |
24 Apr 10:20 Noise increments on Faraday channels
|
new
|
ryanque
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1490 |
24 Apr 11:03 Transmitters turned off
|
new
|
ryanque
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1491 |
24 Apr 11:44 Calibration of ATR and TR
|
new
|
ryanque
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1492 |
24 Apr 15:21 Calibration of ATR and TR
|
new
|
ryanque
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1493 |
24 Apr 16:26 Restart of EW Drift acquisition
|
new
|
ryanque
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1494 |
24 Apr 17:18 Interference in EW Drift and Faraday
|
new
|
ryanque
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1495 |
24 Apr 18:37 Calibration of ATR and TR
|
new
|
ryanque
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1496 |
24 Apr 18:30 Transmitter Power mesearuments
|
new
|
|
Transmitters
|
normal
|
2014_04_ISR+Imaging
|
| #1499 |
25 Apr 08:27 spark gap adjusments
|
new
|
A. Belleza
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1500 |
25 Apr 11:47 Restart of EW_drift acquisition system
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1501 |
25 Apr 12:25 Restart of EW_drift acquisition system
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1502 |
25 Apr 13:31 spark gap adjusments
|
new
|
A. Belleza
|
Spark Gaps
|
normal
|
2014_04_ISR+Imaging
|
| #1504 |
25 Apr 14:15 Change of the attenuation at the B channel
|
new
|
I. Manay
|
Reception
|
normal
|
2014_04_ISR+Imaging
|
| #1505 |
25 Apr 15:34 Restart of EW_drift acquisition system
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_04_ISR+Imaging
|
| #1506 |
25 Apr 15:35 Testing a repaired TR
|
new
|
A. Belleza
|
Reception
|
normal
|
2014_04_ISR+Imaging
|
| #1508 |
05 May 11:37 Log for Imaging+ISR experiment created
|
new
|
ryanque
|
Start experiment
|
normal
|
2014_05 Imaging+ISR
|
| #1509 |
05 May 11:47 Antenna VSWR
|
new
|
ryanque
|
Antennas
|
normal
|
2014_05 Imaging+ISR
|
| #1510 |
05 May 17:15 Start of the experiment
|
new
|
I. Manay
|
Start experiment
|
normal
|
2014_05 Imaging+ISR
|
| #1511 |
05 May 17:30 Restart of Faraday experiment
|
new
|
I. Manay
|
Reception
|
normal
|
2014_05 Imaging+ISR
|
| #1512 |
05 May 17:55 Change of the attenuation
|
new
|
I. Manay
|
Reception
|
normal
|
2014_05 Imaging+ISR
|
| #1513 |
05 May 18:00 Start of the Imaging experiment
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_05 Imaging+ISR
|
| #1514 |
05 May 19:16 Restart of Faraday acquisition
|
new
|
I. Manay
|
Acquisition system
|
normal
|
2014_05 Imaging+ISR
|
| #1515 |
05 May19:00 Transmitter Power measurements
|
new
|
|
Transmitters
|
normal
|
2014_05 Imaging+ISR
|
| #1516 |
05 May 22:30 Calibration ATRs-TRs
|
new
|
J. Lavado
|
Antennas
|
normal
|
2014_05 Imaging+ISR
|
| #1517 |
06 May 00:09 End Imaging
|
new
|
P. Condor
|
Acquisition system
|
normal
|
2014_05 Imaging+ISR
|
| #1518 |
06 May 02:40 Calibration ATRs-TRs
|
new
|
J. Lavado
|
Antennas
|
normal
|
2014_05 Imaging+ISR
|
| #1519 |
06 May 00:00 Figures of rtis, noise,etc
|
new
|
P. Condor
|
Process/Analysis Program
|
normal
|
2014_05 Imaging+ISR
|
| #1520 |
06 May 06:58 Calibration ATRs-TRs
|
new
|
J. Lavado
|
Antennas
|
normal
|
2014_05 Imaging+ISR
|
| #1521 |
06 May 07:00 Transmitter Power mesearuments
|
new
|
|
Transmitters
|
normal
|
2014_05 Imaging+ISR
|
| #1522 |
06 May19:00 Transmitter Power measurements
|
new
|
|
Transmitters
|
normal
|
2014_05 Imaging+ISR
|
| #1523 |
06 Mayo 23:05 Transmitter Power Measurements
|
new
|
|
Transmitters
|
normal
|
2014_05 Imaging+ISR
|
| #1524 |
07 May 01:48 Problems with Power Amplifier 4
|
new
|
AValdez
|
Transmitters
|
normal
|
2014_05 Imaging+ISR
|
| #1525 |
07 May 00:09 End Imaging
|
new
|
AValdez
|
Acquisition system
|
normal
|
2014_05 Imaging+ISR
|
| #1526 |
07 May 01:45 Calibration ATRs-TRs
|
new
|
AValdez
|
Spark Gaps
|
normal
|
2014_05 Imaging+ISR
|